OpenStudy (anonymous):

Calculate the following integral:

OpenStudy (anonymous):

\[\int\limits_{}^{}\sqrt{x^2+1} dx\]

OpenStudy (anonymous):

radical will go if you make the substitution \(x=\tan(\theta)\) although there might be an easier way

OpenStudy (anonymous):

no, probably not you get \[\int sec^3(\theta)d\theta\] and that you can look up

ganeshie8 (ganeshie8):

it looks good to me, how are you saying it is wrong ?

OpenStudy (anonymous):

@satellite73:knew that: \[\sec^2(x)=\tan^2(x)+1\] and he's right but when you substitute you'll see the truth: \[\int\limits_{}^{}\sqrt{x^2+1}[dx]=\int\limits_{}^{}\sqrt{\tan^2(\Theta)+1}[d(\tan(\Theta))]\] look at the dx's we need derivative there

OpenStudy (anonymous):

If x = Tan(Theta) dx = sec^2 (Theta) d(Theta)

ganeshie8 (ganeshie8):

\[\int\limits_{}^{}\sqrt{x^2+1}[dx]=\int\limits_{}^{}\sqrt{\tan^2(\Theta)+1}[d(\tan(\Theta))] \\ = \int \sqrt{\sec^2(\Theta)} \sec^2(\Theta) d\Theta = \int \sec^3\Theta d\Theta \]

ganeshie8 (ganeshie8):

Am I missing something here ?

OpenStudy (anonymous):

Nope. You are right on the money, Ganeshie8. Ifrimpanainte, when we do a substitution, you are correct that we need to substitute in for dx. You got arrived at d(Tan(Theta)), but we can break this down further by actually taking the derivative of Tan(Theta) with respect to Theta. \[\frac{d[Tan(\Theta)]}{d \Theta} = \sec^2(\Theta)\] \[=> d[Tan(\Theta)] = \sec^2(\Theta) d \Theta\]

OpenStudy (anonymous):

You haven't completed it. It's far too early to tell whether Wolfram Alpha agree with us or not (It does)

OpenStudy (anonymous):

Okay now I'm stucked at this point: \[\int\limits_{}^{}\sec^3(\Theta)d(\Theta)\]

OpenStudy (anonymous):

@mathmale